Write an equation of the line given the following the information. Put your answer in slope-intercept form.
Through (7, -9) and perpendicular to 2x + 3y = 9

Found 2 solutions by TimothyLamb, stanbon:
Answer by TimothyLamb(4379)

slope of new line = -1/(-2/3) = 3/2
---
y + 9 = (3/2)(x - 7)
y = (3/2)x - 7(3/2) - 9
y = (3/2)x - 21/2 - 18/2
y = (3/2)x - 39/2
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php

Answer by stanbon(75887)

Write an equation of the line given the following information.
Put your answer in slope-intercept form.
Through (7, -9) and perpendicular to 2x + 3y = 9
---
The given equation has slope = (-2/3)
Your equation must have slope = 3/2
---------
Form of the desired equation:
y = mx + b
Substitute and solve for "b":
-9 = (3/2)7 + b
-18/2 = 21/2 + b
b = -39/2
--------
Equation::
y = (3/2)x - (39/2)
--------------
Cheers,
Stan H.
--------------